OpenStudy (anonymous):

What is the nth derivative of sin^3x?

OpenStudy (anonymous):

what is the first derivative of your expression?

OpenStudy (anonymous):

3/4(cosx- cos3x)

OpenStudy (anonymous):

i want the answer

OpenStudy (jhannybean):

is your problem \(\large \sin^3 (x) \)?

OpenStudy (anonymous):

ya

OpenStudy (jhannybean):

Use the chain rule. \[\large \sin^3 (x) = [\sin(x)]^3\]\[\large \frac{d}{dx}[\sin (x)]^3 * \frac{d}{dx}[\sin(x)]\]

OpenStudy (jhannybean):

Why is this question still open??!

OpenStudy (anonymous):

i want the full solution.

OpenStudy (anonymous):

LOL

OpenStudy (jhannybean):

you mean you havent tried solving it for the past week?!!? cmon...

OpenStudy (anonymous):

Using Bruno's Formula isn't easy. It's a lot of work. But one would use that if you were to actually provide a solution unless there is an easier way.

OpenStudy (callisto):

Maybe diff. this is better?!\[sin(3x) =3sinx - 4sin^3(x)\]\[sin^3(x) = \frac{3sin(x)-sin(3x)}{4}\]

OpenStudy (anonymous):

@Callisto what's that for?

OpenStudy (callisto):

Since \(sin^3(x) = \frac{3sin(x)-sin(3x)}{4}\) Differentiate sin^3(x) is the same as differentiate \(\frac{3sin(x)-sin(3x)}{4}\) To save some work . If you keep differentiate sin^3x, not only need you apply chain rule, you have to apply product rules in finding derivatives later. If you differentiate \(\frac{3sin(x)-sin(3x)}{4}\), all you need is chain rule (for sin(3x)), it becomes easier.

OpenStudy (anonymous):

how's that easier lol. you will be going through sin and cosine as you keep differentiating forever regardless of what you differentiate. In both situations you have to use the chain rule.

OpenStudy (callisto):

But you don't need product rule :D

OpenStudy (callisto):

\[\frac{d}{dx}sin^3x\]\[=3sin^2xcosx\]Then you need product rule when you diff. it again

OpenStudy (anonymous):

i want the nth derivative.can somebody help me wid the solution?

OpenStudy (callisto):

@Sidhulucky Diff. it and observe the pattern.

OpenStudy (anonymous):

@Callisto there is no pattern.

OpenStudy (anonymous):

other than that you will keep hopping between sine and cosine as you take higher derivatives.

OpenStudy (callisto):

That's the pattern

OpenStudy (anonymous):

lol what you show there is basic sine and cosine without any powers. same doesn't go for this case.

OpenStudy (callisto):

That's why we need \[sin^3(x) = \frac{3sin(x)-sin(3x)}{4}\]

OpenStudy (callisto):

If we change \(sin^3x\) into \[\frac{3sin(x)-sin(3x)}{4}\] Then, it is *almost* the same as finding the nth derivative of sin(x)

OpenStudy (anonymous):

mmk. using that "yahoo" way of showing it, it would work.

OpenStudy (callisto):

LOL! Using @.Sam. 's way too :P http://openstudy.com/study#/updates/4f74445ce4b0b478589d9a0c

OpenStudy (jhannybean):

It's basically using the power rule for sin(x).